If 1. on a scale equals 16 oz. Then how would 3 oz. appear on the scale?
Found 2 solutions by nyc_function, stanbon:
Answer by nyc_function(2741)

1/x = 16/3
16x = 3
x = 3/16 or 0.1875 in decimal form.

Answer by stanbon(75887)

If 1. on a scale equals 16 oz. Then how would 3 oz. appear on the scale?
---
Use a proportion:
x/3oz = 1/16oz
---
x = (3/16)
=================
Cheers,
Stan H.
==============